### v1

@@ 4.9 Stretch
The stretching deformation successfully simulated peristaltic locomotion. The forward motion results from wave propagation through contraction and extension of the robot segments along the movement direction (Fig. 5). With bioinspired robotics applications, Nguewou-Hyousse et al. (2019) modeled a caterpillar soft robot using a network of linear and torsional springs and massless rods, modeled as planar discrete elastic rods (Goldberg et al., 2019).
4.10 Discussion on rod models
The analyzed works demonstrate a steady interest in rod models for continuum and soft robots. As shown in Fig. 6, the number of published papers on rod-based models increases.
4.10.1 On the Deformations
Researchers developed numerous rod-based models reformulating and adapting classical rod theories in the soft robotic field. Herein, we reviewed diverse models, which described different combinations of strain modes, giving rise to nine deformation classes (Fig. 4).
The most common class in literature is Bend, with the of the examined papers. This popularity is due to the propensity of slender bodies to bend. In fact, due to the assumption of , with cross-section diameter, the bending stiffness turns out smaller than the stretching or shear ones, easing the bending deformation. In addition, most used actuator paths are linear, which excite predominantly bending modes. These reasons justify the approximation of considering only the bending mode. Furthermore, we can observe that most models in the Bend class leverage the EBRT. However, as summarized in Table 1, several models employed the general CRT. While this represents a computational surplus for trivial bending structures, the generalization of Cosserat models pays off if the model gets deployed in physical interaction tasks.
The next class is All Deformation, with the of the examined papers. As expected, only models based on the CRT achieve all deformations (Table 2). This group of models received a deep study in the last decade, thanks to the rising interest in physical interaction with unstructured environments where all strain modes are significant. Indeed, these models are useful to describe accurately the deformations of slender robots caused by friction (e.g., anisotropic friction for locomotors) or by different mediums (e.g., water). In addition, the most common actuation sources (i.e., pneumatic and cables) excite elongation and compression, causing a significant change in the dimension workspace.
The third class accounting for of the investigated studies is Bend & Twist. These models mainly describe robots in which the unshearability and the unstretchability approximations hold, as summarized in Table 3. Including the twisting deformation allows the representation of backbones with generic 3D curves. However, the twisting deformation can be excited only by specific actuator paths (e.g., helicoidal cables) or asymmetric actuators.
Another significant class, summarized in Table 4, is Bend & Stretch () for the wide diffusion of linear extensible soft pneumatic actuators and tendon-driven robots. Such actuators extend or compress upon solicitation, which induces the continuum body to stretch and bend thanks to their radial disposition on the cross-section.
The remaining five categories are less explored. While the Bend, Twist & Stretch () is a particular case of the All Deformations class, the other classes comprehend few works. The lack of work is due to the difficulty of exciting these specific strain modes with the current actuation technologies or the small number of applications in which the assumptions are valid. In addition, fewer studies that explored shearing per se could be due to the tendency to consider shearing an undesirable or negligible deformation. However, shearing could be significant when there is a prominent environment interaction, as observed in peristaltic locomotion (Hemingway and O’Reilly, 2021) or complex manipulation (Grazioso et al., 2019).
4.10.2 On the Spatial Discretization
One characterizing feature of rod models is the spatial discretization method, which significantly impacts the model’s accuracy and complexity. For simulation and control purposes, the discretized model must be computationally efficient and as accurate as possible. In the analyzed literature, exist two main discretization techniques: Strain Parameterization (Sec. 2.2.3) and Discrete Elastic Rod (DER) (Sec. 2.2.4). The former is based on a discretization directly in the configuration space, allowing the user to choose the degrees of accuracy. In addition, it is possible also to neglect specific strain modes to reduce the computational time. The latter discretizes the length of the rod, considering also the external forces applied to the nodes. Furthermore, most of the existing spatial discretization algorithms are already implemented in dedicated simulators, such as SoRoSim (Mathew et al., 2022) for SP and PyElastica (Naughton et al., 2021) for DER. Straddling geometric and mechanical models, some examined works proposed to fit discrete points with functions, such as PH or Euler’s spirals, which is useful, especially for Shape Estimation.
4.10.3 On the Identification and Validation
to assess the correctness of the rod model,
to estimate the unknown parameters.
In the absence of precise FEM models, lumped-mass models with a high number of DoFs were employed. Otherwise, researchers resort to model optimization to match experimental data. In particular, when the robot morphology is not uniform, researchers usually estimate geometrical properties (e.g., cross-sectional area ) and material characteristics (e.g., elastic modulus , damping coefficient ). Experiments typically conducted are with the robot at different base orientations subject to external loads or at the tip. Most rod models are identified and validated through motion data collected with predefined actuation profiles . However, the literature needs rigorous mechanical tests to yield more accurate models, especially for unstructured physical interactions with the environment. Concerning performance metrics, rod models predicted experimental tip positions with Euclidean errors in the range . However, only a few studies assessed the angular errors, which play a role in dexterous manipulators. In summary, the wide use of experimental data suggests an increasing need to reduce the reality gap to develop more accurate controllers.
4.10.4 On the Impact of Materials and Manufacturing
Manufacturing techniques provided continuum and soft robots with unprecedented dexterity (Wallin et al., 2018). However, manufacturing uncertainties may impact model accuracy. Recent rod models addressed manufacturing uncertainties considering variations in material (Alessi et al., 2023a) and geometric properties (Eugster et al., 2022). Despite these advancements, some gaps persist, particularly in capturing hysteresis effects or integrating the dynamic restoring process of self-healing materials into rod-based models. Hysteresis, a prevalent phenomenon in soft materials, strongly impacts the accuracy and predictability of the robot’s behavior. Furthermore, incorporating self-healing materials can potentially enhance the durability and longevity of soft robots but remains relatively unexplored in rod theories (Terryn et al., 2017). Addressing these challenges could advance the capabilities of continuum and soft robots, enabling their integration into diverse applications and real-world tasks.
5 Controllers using rod models
@@
Controlling continuum robots can be very challenging, due to the intrinsic features of this type of robots. Material nonlinearities (e.g., hysteresis), the infinite-dimensional configuration space, and the under-actuated nature of the system make the direct implementation of well-known controllers for rigid manipulators difficult or impossible.
The literature on the control of rigid robots primarily focuses on controlling the end-effector in terms of pose, velocity, or force. In the realm of continuum robots, the natural equivalent of this task is tip control. To exploit the continuum nature of soft mechanisms and achieve tasks that are impossible for their rigid counterparts, the researchers have proposed two other tasks: Multi-Point (MP) control and shape control. The goal of MP control is to control the pose of a set of cross-sections, while shape control aims to regulate the backbone of the robot, fully utilizing the potential of the continuum structures.
@@
To tackle the inherent properties of continuum robots, rod theories have been widely used for control design, resulting in an increasing trend of works in the last decades, as shown in Fig. 6.
The developed controllers fall into two different categories: model-based and learning-based.
5.1 Model-based Controllers
@@
The Model-Based (MB) control strategy involves leveraging prior knowledge of the robot to guide it toward achieving a specific task. The motivation behind developing MB controllers is the possibility of mathematically guaranteeing the stability and performance of the controlled system. However, implementing controllers in physical prototypes can be challenging, often requiring an extensive identification procedure. This context sees the spatial discretization technique as crucial to balance accuracy and computational efficiency. Therefore, the model selection and the discretization technique are integral parts of the control design process.
5.1.1 Formalization
@@
To formalize the control problem, let us consider a generic configuration vector, whatever discretization algorithm is used. According to Della Santina et al. (2023), the shape control problem can be particularized in two sub-problems: (i) regulation and (ii) trajectory tracking (TT). In the last case, the goal of shape control is to find an input such that
@@
where are the desired configuration and its time derivative, respectively. The condition (48) can be particularized in the regulation case, imposing .
Similarly to rigid robots, it is possible to formalize the tip control problem, rewriting (48) in the task space, such as
@@
where and are the desired pose and velocity twist of the tip. Similarly, the tip pose regulation condition can be obtained, particularizing (49) with . Lastly, the MP control problem can be formalized by extending (49) in a set of cross-sections’ desired poses and velocity twists.
@@
To solve these control problems, it is possible to represent the MB control framework in a general scheme, shown in Fig. 7. The block “Sensors” refers to the sensors mounted on the robotic platform, mapping in the measurements . The “Estimator” block implements the Shape Estimation/Reconstruction algorithms, providing an estimation of , , defined as , . Finally, the MB controller block computes the input , processing the error between the estimated configuration and the desired one. More in-depth, the control input is composed of two actions (Fig. 7). The former is the feedforward action , depending on the chosen model, the estimated state and the desired trajectory. The latter is the feedback action that depends on the error w.r.t. the desired configuration and its time derivative. Finally, the specific control algorithm computes the control output combining the feedforward and feedback terms.
5.1.2 Inverse Kinematics
One of the first approaches proposed in the MB literature is the Inverse Kinematics control. The main idea is to use the inverse kinematic model to find the actuation input such that the tip reaches a desired pose or velocity twist. In the above case, the rod-based kinematic models enable the design of controllers that account for the continuum robots’ deformations.
In the planar case, the EBRT provides a simple and efficient kinematic model. Ataka et al. (2020a) proposed a kinematic pose control for an inflatable eversion CSM for trajectory tracking in a planar environment. Assuming a pure bending robot, they used EBRT to model the effect of internal inflation pressure on each bending segment of the prototype. A further advancement in the modeling is the characterization of the variable structural stiffness caused by changes in inflation pressure. Additionally, an observer based on the Extended Kalman Filter estimates the unknown model parameters online from chamber pressure and bending angle measurements. These parameters approximate the actuator-space Jacobian matrix online, which controls the CSM’s tip pose under various stiffness conditions. Simulation and experimental validation using an inflatable CSM made of unstretchable fabric demonstrated that the proposed method achieved position control in a planar environment despite different robot inflation pressures (i.e., stiffness values).
Mauzé et al. (2021) achieve micrometer positioning accuracy with a PCM made of three planar kinematic chains linked to a rigid moving platform. Each chain has a prismatic actuator and a flexible rod that deforms continuously, with bending and twisting deformations modeled with the CRT. A regular Cartesian controller based on inverse kinematics is implemented using the parameters identified for the forward kinematic model. The controller is validated for a circular and square trajectory. For both, large orientation displacements of the platform increase the angular error.
A closed-loop multi-point orientation control strategy for discretely magnetized continuum manipulators was developed (Richter et al., 2021). A non-homogeneous magnetic field model conforming to Maxwell’s equations was formulated and applied to the actuation system to predict the exerted wrenches on the manipulator. The approach is demonstrated on a manipulator containing two permanent magnets actuated inside a non-homogeneous magnetic field. Here, a quasi-static CRT models the manipulator deformation under external wrenches, while a Jacobian provides the actuation inputs. Experiments in simulation and on a real robot demonstrate that the algorithm independently controlled the orientation of two interconnected magnets in a non-uniform magnetic field. However, the control frequency is limited to Hz due to the computational cost of shape reconstruction and the Jacobian.
Campisano et al. (2021) propose a controller for CSM under tip follower actuation. In particular, the strategy is to couple a quasi-static regime and sensor feedback of the actuation sources. This approach achieves (i) rapid computation of the Cosserat rod model, (ii) actuator nonlinearities compensation, and (iii) oscillation reduction. In conclusion, the closed-loop control algorithm is validated experimentally and compared with an open-loop one, to show the effectiveness of the design process.
Renda et al. (2022) proposed an inverse kinematics controller based on the GVS model, using the actuators-strain mapping in (39). The control algorithm is designed for trajectory tracking of a set of points in . The kinematic model is extended, including a map between the length of threadlike actuators and the configuration of the CSM. This formulation covers the general case of actuators’ routing, such as discontinuous or overlapping routing. The control algorithm is validated through simulations on a CSM with parallel and helical actuators’ routing.
5.1.3 Inverse Dynamics
To exploit the mechanical impedance of continuum robots, the controller requires a knowledge of the dynamics of the elastic rod. Furthermore, the dynamic models allow us to consider the influence of external forces, such as gravity or contacts. Therefore, an inverse dynamic model can be used, compensating undesired nonlinear effects and disturbances.
In the case of shape control, Mbakop et al. (2021a) proposed an inverse dynamics controller for a soft finger. The control algorithm is based on a mix of a geometrical and mechanical approach, combining the PH and the Euler-Bernoulli model. While the former allows the reconstruction of the finger’s shape, the latter maps actuators’ magnitude to the positions of PH control points. They experimentally validated the control law with two prototypes: (i) a planar Fluidic Elastomeric Actuator soft finger and a pneumatic CSM. Applying an external tip load to the CSM proved that combining the PH and EBRT improves accuracy and the computational time, w.r.t. using only the PH model.
The subsequent simulation study proposed a new decentralized Inverse Dynamic controller based on the CRT (Doroudchi and Berman, 2021). The controller considers all strain modes and incorporates the system dynamics inside different gain matrices. Each piece or unit of the soft arm is active and independent from the other pieces. This approach increases the computational efficiency and successfully tracks configurations requiring shear deformations with average normalized root-mean-square errors below .
5.1.4 Feedback Linearization
Continuum robots have an inherent under-actuation, which makes control design more difficult. Feedback Linearization (FL) is a widely used method for controlling under-actuated mechanisms (Spong, 1994). The main idea of this control algorithm is to linearize the system with a feedback control action, compensating all the nonlinear terms. The FL controller can be combined with a linear controller (e.g. Linear Quadratic Regulator), in order to achieve the desired performances.
For a task-space control, Rucker et al. (2022) proposed a controller for continuum robots based on a highly under-actuated planar Kirchhoff rod (i.e., the EBRT). The main idea is to show that a maximal coordinate form of the Kirchhoff model is suitable for control purposes, exploiting the under-actuation of the system. The controller is derived by solving a constrained Lagrangian problem, discretizing the length of the Kirchhoff rod. The control algorithm is based on an input-output Feedback Linearization combined with an Sliding Mode Control (SMC) in an outer loop. An observer based on the constrained model formulation provides control feedback, adding corrective virtual external forces based on the estimation error. Lastly, simulations of an actual two-DoFs soft prototype evaluate the control framework and demonstrate the consistency of the algorithms during tip position trajectory tracking. The numerical examples of the SMC also demonstrate robustness to parameter variations.
Caradonna et al. (2024) presented a 2D soft segment with a revolute joint at its base controlled by a FL controller. By analyzing the equilibrium map of the system’s zero dynamics, they proposed a method to employ the system for the pick-and-place task, accounting for the added mass from the object. They utilize the FL controller with the Curvature Parameterization (27) and Strain Parameterization (11) and compare the two discretization techniques emphasizing numerical robustness, computational cost, and accuracy. The control algorithm and pick-and-place strategy validated through simulations demonstrated consistency.
5.1.5 Energy-Shaping
Another well-known control algorithm designed for under-actuated systems is the Energy-Shaping Control (ESC) (Spong, 1996). The main idea is to find an input such that the system reaches a desired energy state. The rod models are particularly effective in this case, thanks to the Hamiltonian formulation of the elastic rods.
Caasenbrood et al. (2021, 2022) proposed an ESC for CSM for tip pose regulation. To apply the control algorithm, the Authors reformulated the GVS model in a Port-Hamiltonian form, showing the passivity condition of the system and enabling the design of the ESC. The control law is validated through numerical examples, proving the capability of the controlled system to reach the desired tip pose. Furthermore, thanks to the properties of the specific control algorithm, both under-actuation and hyper-redundancy are effectively managed. The controlled system demonstrates smooth convergence to the target, and a pseudo-joint strategy similar to that observed in nature.
Chang et al. (2023) introduced an ESC for a CSM inspired by an octopus tentacle employing the CRT with the DER discretization. They modeled three main types of muscles: transverse, longitudinal, and oblique. The transverse muscles necessitate the DER approach due to the invalid assumption of a rigid cross-section. Leveraging the system’s passivity properties, they designed an ESC by solving an optimal control problem based on Hamiltonian principles. Additionally, they developed an efficient method to design a task-specific equilibrium configuration by solving an optimization problem in the group. The model and ESC are validated on the PyElastica simulator, controlling the CSM for reaching and grasping a static object.
Tiwari and Banavar (2023) proposed a ESC for a 2D flexible single-link manipulator, based on CRT. In particular, the model used is geometrically exact, i.e. the manipulator’s configuration space lies on an infinite-dimensional Lie Group. To implement the controller in the discrete-time domain, the Authors used a lie group variational integrator, derived from the Lagrangian variational principle of the flexible single-link manipulator (Tiwari and Banavar, 2023). The results in silico indicate that the discrete closed-loop system achieves shape regulation during point-to-point rotational maneuvers.
5.1.6 Sliding Mode
An alternative to the above controllers for under-actuated systems is the SMC (Xu and Ümit Özgüner, 2008), due to the inherent robustness. The main idea of this strategy is to control the nonlinear system with a discontinuous control signal, forcing it to a specific curve on the phase space (e.g., a line).
An adaptive SMC based on CRT is proposed by Alqumsan et al. (2019). The controller is designed for planar tip position control, which considers the saturation of the actuation sources. Furthermore, the marginal stability of the proposed adaptive control law is provided. The validation is conducted in simulation, where the PDE of the Cosserat rod are solved by the Generalized- method (Chung and Hulbert, 1993), an implicit solver that assures unconditional stability and numerical dissipation for high-frequency noise.
Subsequently, the high nonlinearity of a large-scale and ultralight two-section inflatable CSM made of TPU-coated Nylon is captured by a dynamic Cosserat rod recursively to predict the behavior of the thin-walled and incomplete continuum structure (Li et al., 2021). An extended inverse dynamic model compensates for the error in a feed-forward manner. Moreover, stiffness and damping matrices are calibrated on deflection data and a weight release experiment, respectively. The study also compares the rod model versus a rigid-body and a PCC model. Finally, three closed-loop model-based controllers for trajectory tracking are compared to further demonstrate the viability of the CRT. Here, an adaptive SMC generated the desired pressure differences in the presence of model uncertainties. The chattering phenomenon caused by the conventional SMC is attenuated by integrating a fuzzy logic control, using continuous membership functions to replace the discontinuous .
Mishra et al. (2023) proposed a Terminal SMC (Venkataraman and Gulati, 1993) for tip position TT control. The model on which the controller is based consists of a combination of CRT and fractional Bounc-Wen model (Mishra et al., 2022), taking into account the hysteresis effect. In this work, the Terminal SMC is robust, fast, non-singular, and with exponential convergence. In addition, an obstacle avoidance algorithm is presented, using a combination of cumulative danger field and damped least square methods. The model and the controller are validated in a simulated version of Robotino-XT (Wang et al., 2022a).
5.1.7 Model Predictive Control
The Model Predictive Control (MPC) (Rawlings et al., 2017) is a well-known controller based on iterative and finite-horizon optimization. Furthermore, it is possible to optimize while satisfying constraints, such as actuators’ saturation.
Towards MIS, Soltani et al. (2017) introduce a 3D, stable, non-linear model-based hybrid force/position controller to apply a constant force to the cardiac tissue while tracking the desired trajectory. A quasi-static Cosserat rod models the distal shaft of tendon-driven catheters, reformulated for controller design. The position controller is based on a non-linear MPC tracker satisfying input constraints and local stability conditions. The position control operates in an inner loop, while the force control is in an outer control loop. Simulations conducted with a bidirectional ablation catheter predict the control performance applied in catheterization. Finally, experimental validation is carried out on a real pig heart, though the authors suggest the necessity of more computational efficiency and incorporation of imaging modalities with position tracking systems.
Recently, a novel motion planning method of fish-like soft robots actuated by macro-fiber composite pairs was developed (Barbosa et al., 2023). The structure should mimic oscillatory and undulation movements, accomplished if the amplitude of the tail motion is greater than that of the head motion. The motion planning adopts MPC for generating the actuation signals. This model-based strategy requires a small-sized model. Therefore, they extracted a linear modal state-space model from a free-free dynamic EBRT considering the electro-mechanical coupling of the actuator pairs. Moreover, design strategies like the addition of concentrated and distributed masses were addressed to mimic fish-like motion, guaranteeing suitable mode shapes for the structure. Four numerical case studies demonstrated the capability of the proposal for deriving bounded input signals that generate oscillatory and undulation movements even with disturbances.
5.1.8 Miscellaneous
An open-loop optimal control method to control an octopus-inspired planar arm modeled with CRT was proposed (Wang et al., 2021b). The main goal is to use the optimal control approach to replicate two stereotypical motions: (i) reaching (i.e., bending propagation from the base to the tip) and (ii) fetching (i.e., creation of pseudo-joints). The functional cost of the optimal problem is the elastic potential and actuation energy of the arm, with a terminal cost related to the specific task. An iterative forward-backward algorithm solves the optimization. The proposal validated in silico demonstrated the emergence of propagating waves, as observed in stereotypical strategies.
Conversely, Doroudchi et al. (2018) designed three controllers based on the dynamics of a linear Euler-Bernoulli rod. Thanks to the linearized form of the EBRT, the authors can apply the optimal control theory easily. The novelty here is that the optimal controllers are decentralized, with a modular soft robotic arm where each piece has independent actuation and sensorization. They observe that ill-conditioned system matrices, common in discretized beam models, may hinder control convergence. The iterative algorithm is also very sensitive to the initial condition. Furthermore, a third controller based on the gradient descent method provides a bounded solution for the controller gains and reduces the sensitivity to initial conditions. Finally, simulations validated the approach.
To control aerial continuum manipulation systems, Hashemi et al. (2023) proposed a terminal synergetic controller to accomplish tip position TT. After presenting a novel platform that integrates a quadrotor with a tendon-bent CTR, they designed a two-loop control architecture. In particular, the first loop implements a terminal synergetic controller to ensure the continuum robot’s tip follows the desired trajectory. The second loop’s controller manages the quadrotor’s position and velocity. In addition, this last control algorithm is based on a combination of synergetic control and a hybrid feedback law. For the first time, a synergetic controller is applied in the literature, showing robustness against disturbances and model uncertainties. Finally, they validate the control framework simulating an aerial continuum manipulation system.

### v2

@@ 6.9 Stretch
The stretching deformation successfully simulated peristaltic locomotion. The forward motion results from wave propagation through contraction and extension of the robot segments along the movement direction. With bioinspired robotics applications, Hyousse2019distributed modeled a caterpillar soft robot using a network of linear and torsional springs and massless rods, modeled as planar discrete elastic rods (goldberg2019planar).
7 Model-based Controllers for Rod Models
@@
Controlling continuum and soft robots can be challenging due to their intrinsic features. Material nonlinearities (e.g., hysteresis), the infinite-dimensional configuration space, and the under-actuated nature of the system make the direct implementation of well-known controllers for rigid manipulators difficult or impossible. The control literature for rigid robots primarily focuses on controlling the end-effector in terms of pose, velocity, or force. For continuum robots, the natural equivalent task is tip control. To exploit the continuum nature of soft mechanisms and achieve tasks that are impossible for their rigid counterparts, the researchers proposed two other tasks:
Multi-Point (MP) control, to regulate the pose of a set of cross-sections,
Shape control, to regulate the robot backbone, fully utilizing the potential of continuum structures.
@@
To tackle the inherent properties of continuum robots, rod theories have been widely used for control design, resulting in an increasing trend of works over the last decade. These controllers generally fall into two categories: model-based and learning-based. This section deals with MB approaches, while the Sec. 8 covers the learning-based.
@@
The MB control strategy involves leveraging prior knowledge of the robot to guide it toward achieving a specific task. The motivation behind developing MB controllers is the possibility of mathematically guaranteeing the stability and performance of the controlled system. However, implementing controllers in physical prototypes can be challenging, often requiring an extensive identification procedure. This context sees the spatial discretization technique as crucial to balance accuracy and computational efficiency. Therefore, the model selection and the discretization technique are integral parts of the control design process.
@@
To formalize the control problem, let us consider a generic configuration vector, whatever discretization algorithm is used. According to della2023model, the shape control problem can be particularized in two sub-problems: (i) regulation and (ii) trajectory tracking (TT). In the last case, the goal of shape control is to find an input such that
@@
where are the desired configuration and its time derivative, respectively. The condition (42) can be particularized in the regulation case, imposing . Similarly to rigid robots, it is possible to formalize the tip control problem, rewriting (42) in the task space, such as
@@
where and are the desired pose and velocity twist of the tip. Similarly, the tip pose regulation condition can be obtained, particularizing (43) with . Lastly, the MP control problem can be formalized by extending (43) in a set of desired poses and velocity twists of the cross sections.
@@
To solve these control problems, it is possible to represent the MB control framework in a general scheme, shown in Figure 5. The block “Sensors” refers to the sensors mounted on the robotic platform, mapping in the measurements . The “Estimator” block implements the Shape Estimation/Reconstruction algorithms, providing an estimation of , , defined as , . Finally, the MB controller block computes the input , processing the error between the estimated configuration and the desired one. Following this general scheme, many works on MB control occurred, thanks to the development of accurate and computationally efficient rod models. Below, we review the contributions, while Table 3 reports the key points.
7.1 Collocated Form
To address the underactuated nature of continuum robots, a recent method proposed by pustina2024input reformulates the dynamic model (14) in a collocated form. This approach decouples the actuated variables from the passive ones via a change of coordinates, . The resulting state vector is partitioned into actuated coordinates, , and passive coordinates, .
A valid change of coordinates can be expressed as
where is the coordinate transformation in the actuated coordinates . For continuum robots modeled with CRT, has a clear physical interpretation. The actuated coordinates correspond to the lengths of the individual actuators, such as
where denotes the length of the -th actuator.
The change of coordinate (45) is significant because it associates a subset of the generalized coordinates with a physically meaningful quantity that can be easily measured (pustina2024input). For example, for a tendon-driven CSM, the actuated variables can be directly obtained using encoders.
In these collocated coordinates, the EoM take the partially decoupled form, such as
where and are the inertial, Coriolis, gravitational, stiffness, and damping matrices, respectively, expressed in the collocated coordinates .
Adopting this change of coordinates is promising for facilitating the transfer of control methods developed for underactuated rigid systems. For instance, pustina2024input demonstrates the application of classical controllers—such as PD with gravity and elasticity compensation (borja2022energy; borja2022role) and its variant P-satI-D (pustina2022p), a PID controller with integral saturation—directly to the collocated variables. Finally, this formulation not only enhances the understanding of the system dynamics and simplifies control design, but also enables the design of linear controllers in the collocated coordinates without approximations. Notably, the control laws remain nonlinear due to the inherent nonlinearity of the mapping .
7.2 Control by Model Inversion
The most direct approach to controlling a continuum robot is through its kinematic model. This strategy, known as Inverse Kinematics (IK), computes the actuation inputs required to drive the robot’s tip to a desired pose. Early studies demonstrated effective planar trajectory tracking and vibration suppression using simplified models such as EBRT (ataka2020model; gravagne2003Large). To better handle real hardware complexities, these models were extended with online observers, which compensate for phenomena such as variable stiffness caused by changes in inflation pressure (ataka2020model).
To achieve higher accuracy and address complex spatial deformations, recent work has shifted towards the more comprehensive CRT. This framework enables advanced capabilities, including micrometer-level positioning in parallel continuum robots (mauze2021micrometer) and multi-point orientation control along a manipulator’s body (richter2021multi). The accuracy of CRT comes at a high computational cost, limiting real-time use. Researchers have therefore turned to quasi-static formulations and actuator nonlinearity compensation to improve efficiency (campisano2021closed). These kinematic models are continuously being generalized to accommodate more complex physical designs, including manipulators with arbitrary, discontinuous, or overlapping actuator routings (renda2022geometrically).
While kinematic models determine the robot’s shape, dynamic models are required for high-speed maneuvers and interactions with the environment. Controllers based on Inverse Dynamics (ID) account for forces such as inertia, gravity, and external contacts. Such methods achieved full shape control, with some approaches combining geometric methods with EBRT to balance accuracy and computational speed (mbakop2021inverse). Because dynamic models are computationally demanding, recent work explored decentralized control architectures that distribute computation across modules, improving real-time feasibility (doroudchi2021configuration).
7.3 Managing Nonlinearities and Under-actuation
A central challenge in continuum robotics arises from the strong nonlinearities and inherent under-actuation of these systems. Feedback Linearization (FL) is a powerful technique that cancels nonlinear terms, transforming the system dynamics into a simpler, linear form. FL was successfully applied in cascade control architectures to compensate for gravity and bending effects, thereby improving tracking performance (falkenhahn2016dynamic). However, its effectiveness depends heavily on model accuracy. To address this limitation, rucker2022task proposed a hybrid approach that combines an inner-loop FL controller with a robust outer-loop strategy such as Sliding Mode Control (SMC), enabling robustness against model uncertainties (rucker2022task). Practical aspects also matter: research shows that the choice of discretization scheme can significantly influence numerical robustness, computational cost, and accuracy (caradonna2024model).
An alternative approach is the Energy-Shaping Control (ESC). Instead of canceling dynamics, ESC reshapes the system’s potential energy such that the desired configuration becomes a stable energetic minimum. By reformulating the dynamics in a Port–Hamiltonian framework, this method can formally guarantee stability and passivity, allowing control of both under-actuated and hyper-redundant systems (caasenbrood2021energy; caasenbrood2022energy). ESC often produces smooth, natural motions that exploit the robot’s intrinsic compliance. It was applied even to complex bio-inspired systems such as octopus tentacles, where optimal control was used to identify shaping laws for advanced models that include non-rigid cross-sections (chang2023energy).
Discrete-time control is necessary for practical implementations, even though most of the ESC literature concentrates on continuous-time models. Structure-preserving numerical techniques, such as Lie group variational integrators, are used to maintain the stability and passivity guarantees in a discretized system (tiwari2023discrete). These integrators allow for efficient stabilization of flexible beams while preserving the geometric structure of the manipulator’s configuration space.
7.4 Robust and Optimal Control
Robust controllers for continuum robots were designed to mitigate model uncertainties. SMC is a widely adopted strategy for ensuring robustness against model uncertainties and external disturbances. Adaptive variants of SMC were developed to handle limitations such as actuator saturation while guaranteeing marginal stability (alqumsan2019robust). A classical drawback of SMC is the high-frequency “chattering” in the control signal. This issue was mitigated by integrating fuzzy logic to smooth the discontinuous control law, leading to smoother actuation profiles (li2021modeling). The methodology continues to evolve with advanced variants, such as terminal SMC, which achieves faster finite-time convergence. It was combined with algorithms for high-level tasks such as obstacle avoidance (mishra2023trajectory).
When performance is critical and constraints must be explicitly respected, Model Predictive Control (MPC) is the method of choice. MPC solves a finite-horizon optimization problem at each time step, enabling it to naturally enforce input bounds while optimizing complex objectives. This capability makes it particularly suitable for medical applications, such as catheter control, where the robot must apply a constant force to tissue while tracking a trajectory (soltani2017soft). MPC is also highly effective for motion planning, as it can generate the complex, oscillatory actuation signals required for behaviors like fish-like swimming, even under disturbances (barbosa2023motion). However, the limitation of MPC remains its computational cost, which often necessitates model simplifications (e.g., linearization) for real-time implementation.
Beyond traditional methods, emerging paradigms include optimal control for bio-inspired motions (wang2021optimal), decentralized robust control for modular architectures (doroudchi2018decentralized), and synergetic control (hashemi2023robust), which has recently shown robustness to disturbances in continuum robots control.